# Seeing the Trees for the Forest: Rethinking Weakly-Supervised Medical Visual Grounding

### v1

@@ 1 Introduction
Visual grounding (VG) is the task of localizing the specific region in an image that corresponds to a given text description. VG is a prominent research area at the vision-language intersection [14, 17, 31], enabling applications like visual question answering  [23, 52], human-robot interaction [47, 51], and data annotation [24, 56]. The ability of linking textual mentioned diseases to specific pathological image regions enhances interpretability [53], aiding radiologists in assessing model decisions. Additionally, VG integrates effectively with large language models (LLMs) to enhance multimodal understanding, e.g., refining 3D object localization through interleaved reasoning and grounding [67], summarizing long text with LLM to improve VG [68], and introducing iterative vision-language interaction with novel tokens for object detection (VG) and text generation [33]. In summary, VG enables transparency, assists localization, and enhances multi-modal interactions on the patch level of the emerging medical multi-modal models [39, 11, 19].
@@ 2 Related Works
Weakly-Supervised Visual Grounding has the same goal as visual grounding but is achieved using only image-level annotations. Prior methods [55, 37, 20, 54, 13] for weakly-supervised visual grounding rely on pre-trained object detectors for Region of Interest (ROI) localization. They typically project text and ROIs into a joint visual-textual embedding space, transforming the grounding task into a retrieval task. In contrast, detector-free methods [45, 44, 2] perform dense localization for a given query phrase without relying on a pre-trained masks proposal model. There are two main approaches in detector-free methods. First, many works [45, 44, 2, 34] generate pseudo-labels, which are expected to approximate the ground truth segmentation, and train models on visual grounding objective. G [45] and g++ [44] employ layer-wise relevance propagation [8] to integrate relevancy map into the cross-modal fusion. Second, other methods [26, 59, 63] optimize some auxiliary tasks on the weakly supervised data. SelfEQ [22] tune model explanation, thus having no detector or pseudo-label generation. Although visual grounding has been well explored for natural images, studies of its applicability in the medical domain is not abundant. Recently, few works [10, 24] perform supervised medical visual grounding that requires pixel-level annotations, which is expensive in the medical domain.
@@ 3.1 Problem Analysis
High norm on BG tokens, hampering the visual grounding capability. The $L_{2}$ norm of patch tokens $\|v\|_{2}$ indicates the activations of the local token $v$. Fig. 4 shows the norm of patch tokens of CheXzero. Tokens in BG regions are frequently activated with high values. Unlike natural images, medical images are highly canonical, with significant redundancy in BG areas. Hence, high-norm tokens often occur in these healthy BG regions rather than in the disease-associated FG areas. We then show that the high activation of BG tokens hampers the visual grounding. Fig. 4 further compares the norm map and the visual grounding of CheXzero [50]. The model mistakenly grounds on background tokens with high norms.
To quantify how spurious norm activation in the background (BG) regions affects visual grounding (VG), we calculate the $L_{2}$ norm of the tokens to form a norm map. We then evaluate the correlation between these norm maps and VG predictions. Fig. 5 shows two findings. First, VLMs show high norm activations in the background regions, evident by the high Dice score between the response map and the BG regions. Second, the Dice score between VG outputs on the norm maps is significantly higher than the Dice with the actual ground truth (GT). This shows that the models tend to ground on the highly activated regions of the norm map. These findings suggest that high norm activations in BG regions can divert the model away from disease areas, weakening VG accuracy.
@@
Global image [IMG] token is non-representative of the local foreground disease tokens. Current VLMs mostly operate on the global image [IMG] when aligning with the text token. Are they representative of local patch tokens $v_{i}$? This study shows a negative finding. Fig.6 investigates the intra-modal alignment between the global [IMG] token and foreground (FG) and background (BG) tokens using cosine similarity. In the top row, the [IMG] token aligns more closely with BG tokens than with FG tokens. This is further substantiated by the average cosine similarity between the global [IMG] token and individual patches (see Fig. 7).
@@ 4 Experimental settings
Implementation Details. Our experiments are conducted on 4 NVIDIA RTX A100 GPUs. Input images are resized to 224$\times$224 and augmented with RandAugment [12]. We set up an Adam optimizer [16] with a learning rate of 0.008 and a batch size of 512 across all experiments. We cap the maximum length of each text to 77 as suggested by [50]. Our hyperparameter values and schedules were determined empirically on a small validation subset.
@@ 6 Conclusion
This paper provides a diagnostic view of local alignment of medical vision-language models, hampering the weakly-supervised visual grounding. We show that (i) these models assign high norm values to non-disease regions, misdirecting attention, and (ii) intra-modal misalignment between global and local tokens weakens disease-specific grounding. As such, we propose a simple-yet-effective disease-aware prompting (DAP) technique that strengthens disease-region representations while suppressing background interference. This paper constructs a comprehensive weakly-supervised medical visual grounding on three datasets to catalyse the research. On this new benchmark, ours outperforms state-of-the-art methods by up to 20.74% across 3 datasets.

### v2

@@ 1 Introduction
Visual grounding is a prominent research area at the vision-language intersection [15, 18, 33], enabling applications like visual question answering  [24, 55], human-robot interaction [50, 54], and data annotation [26, 59]. Beside concept-based approaches [25], the ability of linking textual mentioned diseases to specific pathological image regions enhances interpretability [56], aiding radiologists in assessing model decisions. Additionally, VG integrates effectively with large language models (LLMs) to enhance multimodal understanding, e.g., refining 3D object localization through interleaved reasoning and grounding [68], summarizing long text with LLM to improve VG [69], and introducing iterative vision-language interaction with novel tokens for object detection (VG) and text generation [35]. In summary, VG enables transparency, assists localization, and enhances multi-modal interactions on the patch level of the emerging medical multi-modal models [41, 12, 20].
@@ 2 Related Works
Weakly-Supervised Visual Grounding aims to localize target objects mentioned in the text without segmentation labels. Detector-based methods [40, 58, 39, 21, 57, 14] adopts two-stage pipeline, using pre-trained object detectors to extract region of interests (RoIs), and then matching texts with the cropped RoI. A recent work [40] generates pseudo-captions for each RoI using BLIP, and matches between the text query and the pseudo-caption. RefCLIP [31], and QueryMatch [10] propose a one-stage pipeline, directly matching the text with the anchor point in the feature map or the DETR query. A state-of-the-art grounding multimodal LLM [42] uses external segmentation models to train LLMs for grounding. Unfortunately, the reliability of medical-target detectors is not comparable with the general domain. In contrast, detector-free methods [48, 47, 2] leverage pseudo labels generated by the explainability map [8].G [48] and g++ [47] re-calibrate visual feature using the image-text similarity map. They also force the model to mask the background region to yield maximum similarity with the text. G++ [47] aggregates the self-attention map having high overlap with the grounding to create a more comprehensive pseudo-label. SelfEQ [23] encourages consistent visual grounding map with diverse text prompting. Existing WSVG methods rely on strong initial grounding map. Yet, current medical vision language models show low image-and-text alignment.
@@ 3.1 Problem Analysis
High norm on BG tokens, hampering the visual grounding capability. The $L_{2}$ norm of patch tokens $\|v\|_{2}$ indicates the activations of the local token $v$. Fig. 4 shows the norm of patch tokens of CheXzero. Tokens in BG regions are frequently activated with high values. Unlike natural images, medical images are highly canonical, with significant redundancy in BG areas. Hence, high-norm tokens often occur in these healthy BG regions rather than in the disease-associated FG areas. We then show that the high activation of BG tokens hampers the visual grounding. Fig. 4 further compares the norm map and the visual grounding of CheXzero [53]. The model tends to ground on background regions with high norm.
To quantify how spurious norm activation in the background (BG) regions affects visual grounding (VG), we calculate the $L_{2}$ norm of the tokens to form a norm map. We then evaluate the correlation between these norm maps and VG predictions. Fig. 5 shows two findings. First, VLMs show high norm activations in the background regions, evident by the high Dice score between the response map and the BG regions. Second, the Dice score between VG outputs on the norm maps is significantly higher than the Dice with the actual ground truth (GT). This shows that the models are inclined to ground on the highly activated regions of the norm map. These findings suggest that high norm activations in BG regions can divert the model away from disease areas, weakening VG accuracy.
@@
Global image [IMG] token is non-representative of the local foreground disease tokens. Current VLMs mostly operate on the global image [IMG] token when aligning with the text token. Is [IMG] representative of local patch tokens $v_{i}$? This study shows a negative finding. Fig.6 investigates the intra-modal alignment between the global [IMG] token and foreground (FG) and background (BG) tokens using cosine similarity. In the top row, the [IMG] token aligns more closely with BG tokens than with FG tokens. This is further substantiated by the average cosine similarity between the global [IMG] token and individual patches (see Fig. 7).
@@ 4 Experimental settings
Implementation Details. Our experiments are conducted on 4 NVIDIA RTX A100 GPUs. Input images are resized to $\times224$ and augmented with RandAugment [13]. We set up an Adam optimizer [17] with a learning rate of 0.008 and a batch size of 512 across all experiments. We cap the maximum length of each text to 77 as suggested by [53]. Our hyperparameter values and schedules were determined empirically on a small validation subset.
@@ 6 Conclusion
This paper provides a diagnostic view of local alignment of medical VLM, hampering the weakly-supervised visual grounding. We show that (i) these models assign high norm values to non-disease regions, misdirecting attention, and (ii) intra-modal misalignment between global and local tokens weakens disease-specific grounding. As such, we propose DAP that strengthens disease-region representations while suppressing background interference. This paper constructs a comprehensive weakly-supervised medical VG on three datasets to catalyze the research, where DAP ours outperforms state-of-the-art methods by up to 20.74%.
Appendix A Self-enhancement with DAP
The interpretability map $\Phi$ shows a reasonable localization capability as it achieves 33.6% average dice score on MS-CXR. Furthermore, $\Phi$ helps enhance the model by narrowing down the pathological search space for VG. By dampening the influence of background, the model can discover finer pathological signals, which get dominated without the initial interpretability map prompting. Fig. 13 shows that our model further refines the localization ofter $\Phi$ narrows down the search space.
Figure 13: Results of interpretability map, our DAP, and others.
Fig. 14 (Left) plots the Dice score of our model against $\Phi$ on RSNA. Most points lie above the red line, showing our model imprives Dice upon the interpretability map, indicating self-enhancement. Fig. 14 (Right) shows DAP surpasses baselines even when $\Phi$ fails (Dice << 0.3).
Figure 14: Left: Relationship between the performance of the interpretability map and our DAP. Points above the red line (highlighted in orange) indicate self-enhancement cases, where our DAP further improves the Dice score of interpretability maps. Right: Avg. Dice on samples grouped by the quality of the interpretability map. Dice of other VG are shown for comparisons.
Appendix B Hyperparameters tuning
In this section, we study the impact of different hyperparameter settings, including the batch size, learning rate, the weights of loss objectives, text prompts variety and prompt depth. We conduct the hyperparameters tuning on 20% of the MS-CXR [6] dataset.
B.1 Batch Size and Learning Rate
We vary the batch size of $\{128,256,512\}$ and report the result in Table. 5. It is shown that larger batch size has a positive impact on the overall performance.
Table 5: Batch size versus performance.
[…]
128 1.027 0.443 0.340
256 1.038 0.440 0.347
512 1.042 0.449 0.350
We vary the learning rate of $\{1e^{-1},1e^{-2},1e^{-3}\}$ and report the result in Table. 6. The optimal learning rate is $1e^{-3}$. We set batch size to 512 and learning rate to $1e^{-3}$ for other experiments.
Table 6: Learning Rate versus performance.
[…]
$1e^{-1}$ 0.917 0.358 0.288
$1e^{-2}$ 1.054 0.425 0.345
$1e^{-3}$ 1.042 0.449 0.350
B.2 Loss weights
We vary the weights of the Disease-aware Global contrastive loss $\mathcal{L}_{glb}$ and the local contrastive loss $\mathcal{L}_{lcl}$ to evaluate their significance. We set their weights to $\{0.1,1,2\}$ and report the result in Table. 7. It is demonstrated that setting the weights for $\mathcal{L}_{glb}$ to 1 and $\mathcal{L}_{lcl}$ to 0.1 achieves the best overall performance, yielding the highest dice score of 0.350, PG of 0.449, and CNR of 1.042. Notably, increasing the weight for $\mathcal{L}_{lcl}$ to 1 or 2 leads to a reduction in the dice score and slight degradation in PG and CNR, suggesting that overemphasizing localization introduces diminishing returns. $\mathcal{L}_{glb}$ significantly influences segmentation performance, while localization loss has a lesser impact.
Table 7: Loss weights versus performance.
[…]
1 0.1 1.042 0.449 0.350
1 1 1.030 0.445 0.343
1 2 1.036 0.438 0.344
2 0.1 1.036 0.443 0.341
2 1 1.042 0.450 0.344
0.1 1 1.038 0.436 0.343
0.1 2 1.041 0.446 0.345
B.3 Robustness with other interpretability methods.
Chefer et al.’s [8] method was chosen for its strong track record in VG tasks and alignment with the bi-modal nature of VLMs, which is often used in previous VG works in VPT, g, and g++. To assess robustness, we implemented DAP with different methods on RSNA and COVID datasets with results in Tab. 8. It shows that GradCAM and SmoothGrad closely match the original performance, while LRP achieves similar scores to Self-EQ (CVPR2024). As such, DAP generalizes to other interpretability methods but we empirically found that  [8] gives optimal performance.
Table 8: CNR scores of different interpretability methods.
Chefer et al[8] GradCAM SmoothGrad LRP Self-EQ
RSNA 1.630 1.462 1.507 0.973 1.075
COVID 1.018 0.903 0.891 0.612 0.659
B.4 Prompt depth
We investigate the impact of prompting at different layers within the vision encoder. Specifically, we evaluate prompting at the pixel space of the original image, the first half of the encoder, the full encoder, the second half, and exclusively at the last layer. We report the result in Table. 9. The ”last layer” setup achieves the best trade-off between segmentation accuracy and robustness, offering a strong balance between dice and CNR. In contrast, configurations like ”full” or ”last half” slightly enhance CNR but compromise dice, underscoring the effectiveness of focusing on deeper features for balanced performance. In contrast, the ”first only” and ”first half” configurations underperform.
Table 9: Disease aware prompting layer depth versus performance.
[…]
first layer 1.037 0.443 0.341
first half 1.039 0.442 0.343
full 1.043 0.450 0.343
last half 1.046 0.443 0.344
last layer 1.042 0.449 0.350
B.5 Fixed versus Varied text Prompts
We investigate textual prompting strategies by exploring one fixed prompt, and multiple paraphrases per disease class. The settings range from 1 prompt per class to 20 and 50 prompts per class, with results summarized in Table. 10. The findings demonstrate that increasing the number of prompts per class enhances model performance, likely due to expanded vocabulary exposure, which improves the robustness of the text model.
Table 10: Number of text prompts per class versus performance.
[…]
1 0.991 0.417 0.334
10 1.037 0.443 0.344
20 1.036 0.448 0.349
50 1.042 0.449 0.350
Appendix C Few-shot Supervised finetuning performance
We further evaluate the proposed approach in few-shot settings, starting with weakly-supervised training followed by 20-shot fine-tuning using ground truth dense labels. As shown in Table. 11, our proposed DAP achieves results under weakly-supervised settings that are on par with the 20-shot fine-tuned performance of competing methods, demonstrating its efficiency in learning meaningful representations with minimal supervision for visual grounding.
[…]
Method Venue weak 20-shot weak 20-shot weak 20-shot
[48] NIPS22 1.461 1.606 0.284 0.782 0.450 0.512
g++ [47] CVPR23 1.350 1.625 0.467 0.733 0.445 0.456
VPT [36] ICASSP24 1.173 1.599 0.612 0.785 0.468 0.535
DAP Proposed 1.630 1.741 0.747 0.816 0.474 0.551
Table 11: Weakly-supervised and Few-shots finetuning performance on RSNA [49] dataset.
Appendix D Text prompts construction
We utilize GPT-4o [43] to generate disease-centric descriptions for chest X-ray findings. The model is instructed to produce 50 distinct prompts for each disease, explicitly avoiding anatomical location details. The prompt we used is:
For the disease <disease name>, provide a sentence to describe it on chest X-rays. Exclude any reference to anatomical locations and ensure findings are concise, medically accurate, and reflect a professional radiology reporting style.
Appendix E Robustness to flawed $\Phi$
We inject noise into $\Phi$ and investigate how DAP degrades as noise increases. We first consider pixels of $\Phi$ with value >0.3>0.3 as important. Then, we flip top-$k\in\{10,30,50,70\}\%$ of important pixels to 0, and retrain the model to find how good should $\Phi$ be to benefit DAP. We conducted experiments on RSNA and plot the result in Fig. 15. DAP’s performance is indeed correlated with the quality of $\Phi$, which is lower than g++ when $\text{dice}(\Phi,\text{GT})\leq 0.3$ at $k=50\%$. Yet, we note that this is artificial scenario. In practice, $\Phi$ exhibits a reasonable localization capability with 0.340.34/0.420.42/$0.33$ dice score on MS-CXR/RSNA/Covid.
Figure 15: Dice score (left) and CNR (right) of DAP and $\Phi$ against the ground truth of RSNA dataset under different noise levels, compared to the strong baseline g++[44].
Appendix F Qualitative results
We present more qualitative results of our proposed DAP in Fig. 16 and Fig. 17.
Figure 16: Qualitative results of DAP on MS-CXR [6] dataset.
Figure 17: More qualitative results of DAP on MS-CXR [6] dataset.